OpenStudy (anonymous):

1) sin(lnx)dx , from 1 to e

OpenStudy (anonymous):

This is your problem?\[\int\limits_{1}^{e}\sin(lnx)dx\]

OpenStudy (anonymous):

@Amum

OpenStudy (anonymous):

yh

OpenStudy (anonymous):

I would recommend using substitution.

OpenStudy (anonymous):

Substitute \[u=ln(x) =>du=\frac{1}{x}dx\] Find new bounds, and solve it from there.

Parth (parthkohli):

for these bullcrap functions all I can think of is IBP

OpenStudy (anonymous):

When you make the substitution you get: \[e^usin(u)\] To integrate that you would need to use the formular:\[\int\limits_{ }^{ }\exp (\alpha u)\sin(\beta u)du=\frac{ \exp(\alpha u)(-\beta \cos(\beta u)+\alpha \sin(\beta u)) }{ \alpha^2+\beta^2 }\]

OpenStudy (anonymous):

Ok got it, thx